OpenStudy (anonymous):

How can the Laplace transformation L{f(ct)} (where c is a constant) be derived using the definition of a Laplace transformation? (As opposed to finding it by using a table of known Laplace transforms?)

OpenStudy (wmj259):

Calculus 3?

OpenStudy (anonymous):

DE

OpenStudy (amistre64):

whats the definition of a laplace transform?

OpenStudy (amistre64):

\[L\{f(ct)\}=\int^{\infty}_{0}f(ct)~e^{-st}~dt\]

OpenStudy (amistre64):

do you know how to work a by parts ... using a table?

OpenStudy (anonymous):

yes I do I have that the answer is (1/c)F(s/c) I just have to show how to get there and that's what I'm confused on

OpenStudy (amistre64):

if we work the by parts, im sure we will get to that

OpenStudy (anonymous):

okay so if i start the integral should i do integration by parts?

OpenStudy (anonymous):

u= f(ct) ; du=c*f'(ct) ; dv=e^-st dt ; v=(-1/s)e^-st

OpenStudy (amistre64):

\[ \Large \begin{array}c & & \int dt\\(+/-) & d/dt & e^{-st}\\+ & f(ct) & -\frac{1}{s}e^{-st}\\- & c f'(ct) & \frac{1}{s^2}e^{-st}\\+ & c^2 f''(ct) & -\frac{1}{s^3}e^{-st}\\- & c^3 f'''c(t) & \frac{1}{s^4}e^{-st}\\\end{array} \]

OpenStudy (amistre64):

c got a little out of the function at the end lol

OpenStudy (amistre64):

integration by parts gives us a pattern that gets established right?

OpenStudy (anonymous):

yeah i see that pattern, i'm just not sure how the answer in the book got (1/c) on the outside of the inverse laplace transform of (s/c) ?

OpenStudy (anonymous):

what you're saying totally makes sense cause the power of c is increasing so i'm not really sure?

OpenStudy (amistre64):

\[\int_{0}^{\infty}f(ct)e^{-st}~dt=-\underbrace{\sum_{n=0}^{\infty}c^nf^{(n)}(c\infty)\frac{e^{-s\infty}}{s^{n+1}}}_{=0}+\sum_{n=0}^{\infty}c^nf^{(n)}(c0)\frac{e^{0}}{s^{n+1}}\]

OpenStudy (amistre64):

hmm, the other option is to derivative e^(-st) and integrate f(ct) but that might be complicated ....

OpenStudy (anonymous):

okay so the second summation looks very similar to the t^n la place transformation correct?

OpenStudy (amistre64):

yes

OpenStudy (amistre64):

now, are you sure you typed the correct results?

OpenStudy (anonymous):

i'm trying to use the equation drawer but its not working let me see if i can get it to work just to double check

OpenStudy (amistre64):

equation button is not that well equiped. knowing latex allows you to type it in directly to this reply box and avoids the hunt and peck

OpenStudy (anonymous):

\[inverse laplace = f(ct) ; F(s) [laplace] = \frac{ 1 }{ c }F(\frac{ s }{ c }) , c > 0\]

OpenStudy (amistre64):

.. yeah, thats what i thought you typed :) now working it according to the definition

OpenStudy (anonymous):

okay so we're on the right track :)

OpenStudy (anonymous):

so for my answer i would just write the integral above equals the summation and that should satisfy the proof?

OpenStudy (amistre64):

int u dv = uv - int v du let u = e^(-st), du = -s e^(-st) dt let v = 1/c F(ct) + K ; dv = f(ct) sorry, really really had to go ... do something

OpenStudy (anonymous):

haha that okay thanks a lot for your help!

OpenStudy (anonymous):

ohhhhh okay now i see the 1/c can't believe i didn't see that before

OpenStudy (amistre64):

\[\int_{0}^{\infty}f(ct)e^{-st}~dt=\left(\frac 1c F(ct)+K\right)e^{-st}|^{\infty}_{0}+\frac sc\int_{0}^{\infty} \left(F(ct)+K\right) e^{-st}~dt\] trying to code it really lags my computer

OpenStudy (amistre64):

i think K is immaterial here since its a definant integral

OpenStudy (anonymous):

I didn't even think of making the exponential function u ,okay perfect I got it!

OpenStudy (anonymous):

got any time for another laplace problem? ;)

OpenStudy (amistre64):

you got it? good, becuase between trying to code and keep up with mathing it ... i tend to get lost inbetween lol

OpenStudy (amistre64):

maybe, whats teh problem?

OpenStudy (anonymous):

y'' + y = [t, <= t < 1 ] & [0, 1 <= t < infinity] y(0)=0 y'(0)=0 ?

OpenStudy (amistre64):

is that a heaviside? pieced function?

OpenStudy (anonymous):

yes piecewise defined

OpenStudy (anonymous):

i have some of it worked out i'm going to try to upload a picture of it ?

OpenStudy (amistre64):

y'' + y = t, 0 < t < 1 y'' + y = 0, 1 < t < inf

OpenStudy (amistre64):

L{y''} + L{y} = L{t} L{y''} + L{y} = L{0} or are you sposed to do definition and not table?

OpenStudy (anonymous):

on the first one it is 0 <= t < 1 1 <= t < infinity

OpenStudy (anonymous):

i can use the table

OpenStudy (amistre64):

your work is looking fine, but the thing is, the interval is not from 0 to inf sooo, you might have to split it \[\int_{0}^{inf}f(t)e^{-st}dt=\int_{0}^{1}f(t)e^{-st}dt+\int_{1}^{inf}f(t)e^{-st}dt\]

OpenStudy (anonymous):

yeah you're right. so my answer of y = t - sin(t) would only solve the first part of the piece wise function ? or is that part wrong ?

OpenStudy (amistre64):

it would solve it IF there wasnt a change in function, instead of e^(inf) in the normal operations of things, we stop at e^(-s) which doesnt zero out

OpenStudy (anonymous):

alright i'll start from the beginning

OpenStudy (amistre64):

use the definiton and youll see how it alters the setup

OpenStudy (anonymous):

i'm hopeless lol

OpenStudy (amistre64):

L{y''(t)}, [0,1] int[0,1] y''(t) e^(-st) dt u = e^(-st) v = y'(t) dt du = -s e^(-st) dv = y''(t) dt int[0,1] y''(t) e^(-st) dt = [y'(1) e^(-s)- y'(0)] +s int[0,1] y' e^(-st) dt L{y''(t)}, [0,1] = y'(1) e^(-s) +s L{y''(t)}, [0,1] does this make sense?

OpenStudy (amistre64):

y' on the end, got a little '''' happy lol

OpenStudy (anonymous):

okay about to upload my work

OpenStudy (anonymous):

this thing needs emojis or something lol

OpenStudy (anonymous):

really close just a little off

OpenStudy (anonymous):

sorry i was having a hard time using the integral way because my teacher never taught it that way

OpenStudy (amistre64):

im trying to paper it too ... let review, mainly for my sake :) \[\left.L\{y'(t)\}\right|_{0}^{a}=\int_{0}^{a}y'(t)e^{-st}dt=\\ ~~~~~~~~\left[y(a)e^{-sa}-y(0)e^{-s(0)}\right]+s\int_{0}^{a}y(t)e^{-st}dt\] \[\left.L\{y'(t)\}\right|_{0}^{a}=y(a)e^{-sa}-y(0)+sL\{y(t)\}|_{0}^{a} \]

OpenStudy (amistre64):

as a approaches infinity, this is the usual laplace that we are familiar with yes?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

becasue we have the function split up along intervals, then this is a process that we can do.

OpenStudy (amistre64):

a to infnity just gives us y(inf) e^{-inf}-y(a) e^{-sa} right?

OpenStudy (anonymous):

right

OpenStudy (amistre64):

not sure what im getting at here, i just know what the laplace is altered when we dont do a complete interval from 0 to 1 :)

OpenStudy (anonymous):

hmm me either but i definitely see the direction you're going in

OpenStudy (anonymous):

i have to go :( but i think if i spend some more time on it i can get it, thank you very much for all your help!

OpenStudy (amistre64):

ill continue to work at it just becasue i like the question :) have fun

OpenStudy (anonymous):

okay thank you!!

OpenStudy (anonymous):

@amistre64 A more straightforward approach to establish the transform of \(f(ct)\) might be to substitution \(u=ct\), then you get \[\begin{align*}\mathcal{L}\{f(ct)\}&=\int_0^\infty f(ct)~e^{-st}\,dt\\\\ &=\frac{1}{c}\int_0^\infty f(u)~e^{-su/c}\,du\\\\ &=\frac{1}{c}F\left(\frac{s}{c}\right)\end{align*}\] where the last line is recognizing that the \(s\) argument in the integral is replaced with \(\dfrac{s}{c}\).

OpenStudy (amistre64):

yeah, i was a bit stymied at how we got the s/c in there :) the change of variable makes it more sensible to me know.